trains A and B are traveling in the same direction on the parallel tracks. train A is traveling 100 miles per hour and train B is traveling at 110miles per hour. Train A passed a station at 7:10pm if train B passes the same station at 7:22pm. At what time till train B catch up with train A?

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(57967)

trains A and B are traveling in the same direction on the parallel tracks. train A is traveling 100 miles per hour and train B is traveling at 110miles per hour. Train A passed a station at 7:10pm if train B passes the same station at 7:22pm. At what time till train B catch up with train A?
-------------
Train A DATA:
rate = 100 mph ; time = x hrs. ; distance = rt = 100x miles
-----
Train B DATA:
rate = 110 mph ; time = (x-(22/60)) hrs ; distance = 110*(x-(11/30)) miles
--------------------------
Equation:
distance = distance
100x = 110(30x-11)/30)
(10/11)x = (30x-11)/30
Cross-multiply:
300x = 11(30x-11)
300x = 330x - 121
30x = 121
x = 121/30 hr (time for A Train)
----
Time = 7:10 + (121/30) hr
Time = 7:10 + 4 hrs + 0.033333*60 min
Time = 11:10 + 2 minutes
Time = 11:12 pm
======================
Cheers,
Stan H.

Answer by Alan3354(31519)

trains A and B are traveling in the same direction on the parallel tracks. train A is traveling 100 miles per hour and train B is traveling at 110miles per hour. Train A passed a station at 7:10pm if train B passes the same station at 7:22pm. At what time till train B catch up with train A?
-----------------
722 - 710 = 12 minutes = 1/5 hour
Train A goes 20 miles in that time (100 mph * 1/5 hour) and it 20 miles ahead of train B at 722.
Train B is gaining on train A at 10 mph (110 - 100).
20 miles/10 mph = 2 hours.
722 + 200 = 922 PM